# Batalin-Fradkin-Vilkovisky Quantization of FLPR model

### v1

@@
Abstract: We quantize the Friedberg-Lee-Pang-Ren (FLPR) model within the framework of Batalin-Fradkin-Vilkovisky (BFV) formalism. We construct the nilpotent Becchi-Rouet-Stora-Tyutin (BRST) charges using constraints and the fermionic gauge-fixing function by means of admissible gauge conditions. We also derive the BRST invariant effective action and corresponding symmetries of the model in both polar and Cartesian coordinates. Additionally, with the help of BRST charge, we demonstrate that the physical states of the system are annihilated by the first-class constraints, which is consistent with the Dirac formalism.
PACS Nos: 11.15.-q, 11.10.Ef, 11.30.-j
Keywords: BFV quantization; FLPR model; Constrained systems; BRST symmetries.
@@ 1 Introduction
Constrained theories are the cornerstone of modern theoretical physics, as they include gauge theories, gravity theories, super-symmetric theories etc. The conventional methods of quantization are not directly applicable to such theories. Dirac developed a quantization technique to deal with such constrained systems - known as the Dirac formalism [1]. Later, a geometrically motivated approach based on the symplectic structure of the phase space - the Faddeev-Jackiw (FJ) formalism - was proposed [2]. The FJ formalism is also quite popular and widely used (see e.g. [3, 4] and references therein). As far as quantization of such constrained theories is concerned, the Becchi-Rouet-Stora-Tyutin (BRST) formalism is one of the most broadly used techniques [5, 6].
@@
However, an another versatile method for the quantization of constrained system is introduced by Batalin, Fradkin and Vilkovisky and popularly known as the BFV formalism [7, 8]. It provides a systematic Hamiltonian framework for quantizing theories with first-class constraints which also preserves the BRST symmetries of the system. The BFV formalism has been generalized to include second-class theories by employing a conversion technique called the BFT formalism [9]. These methods have been explored in various classical and field theoretical models such as Chiral Schwinger model, Proca theory, QED and $(2+1)$-dimensional Hořava theory [10, 11, 12, 13].
@@
The Friedberg-Lee-Pang-Ren (FLPR) model is a soluble gauge model that describes the motion of a single non-relativistic particle of unit mass governed by a central potential [14]. This is a $(0+1)$-dimensional quantum mechanical model with the characteristics of Gribov ambiguity, inspired by the dimensional reduction of Yang-Mills theory [15]. The BRST invariant path integral for the model is formulated by summing over all Gribov-type copies [16]. In addition, the FLPR model is being quantized using an admissible gauge-fixing condition within the framework of FJ quantization [17]. The system is also explored using a physical projector operator which avoids the gauge-fixing procedure [18]. Moreover, the existence of some novel off-shell nilpotent and absolutely anti-commuting symmetries of the FLPR model is established in [15]. Recently, this model is being investigated by means of supervariable and (anti-)chiral supervariable approaches [19, 21, 20].
@@
The prime motive of our present endeavor is to quantize the FLPR model using an admissible gauge-fixing condition, in both polar and Cartesian coordinates, by employing BFV formalism. We intend to construct the nilpotent and conserved BRST charge and subsequently derive the BRST invariant gauge-fixed effective action and corresponding symmetries. In addition, we plan to show that the physical states of the theory are annihilated by the conserved BRST charge which is in accordance with the Dirac-quantization conditions.
@@
The content of the present manuscript is organized as follows. In section 2, we provide the constraint structure of the FLPR model in polar coordinates using Dirac’s prescription of constraint analysis. The BFV quantization of the model is carried out in subsection 2.1. Our section 3 deals with the constraint analysis of the FLPR model in Cartesian coordinates whereas BFV formulation, in Cartesian coordinates, is presented in subsection 3.1. The section 4 discusses about conserved and nilpotent BRST charges and physicality criteria. Finally, we conclude our results in section 5.
2 FLPR model in polar coordinates: Constraint structure
@@
where $\dot{r}$, $\dot{\theta}$ and $\dot{z}$ are generalized velocities corresponding to the generalized coordinates $r$, $\theta$ and $z$, respectively. Here, $\zeta$ represents a gauge variable and $g$ is coupling constant. The canonical momenta corresponding to the variables $r$, $\theta$, $z$ and $\zeta$ are given, respectively, as
$p_{r}=\dot{r},\qquad p_{\theta}=r^{2}(\dot{\theta}-g\zeta),\qquad p_{z}=\dot{z}-\zeta,\qquad p_{\zeta}=0.$ (2)
@@
${H}_{c}=\frac{1}{2}p_{r}^{2}+\frac{1}{2r^{2}}p_{\theta}^{2}+\frac{1}{2}p_{z}^{2}+\zeta(gp_{\theta}+p_{z})+V(r).$ (3)
@@
There are no further constraints in the theory. Thus, the theory possess only two constraints, which are first-class in nature [1, 17].
@@ 2.1 BFV quantization
In order to make use of the BFV formalism, we introduce two sets of canonical ghost and anti-ghost variables $\big{(}{\cal{C}}^{i},{\cal\bar{P}}_{i}\big{)}$ and $\big{(}{\cal{P}}^{i},{\cal\bar{C}}_{i}\big{)}$ along with the auxiliary variables $\big{(}{N}^{i},{B}_{i}\big{)}$, such that they satisfy following super-Poisson bracket algebra
$\begin{gathered}\{{\cal{C}}^{i}(x),{\cal\bar{P}}_{j}(y)\}=\delta_{j}^{i}\delta(x-y),\\ \{{\cal{P}}^{i}(x),{\cal\bar{C}}_{j}(y)\}=\delta_{j}^{i}\delta(x-y),\\ \{{N}^{i}(x),{B}_{j}(y)\}=\delta_{j}^{i}\delta(x-y),\end{gathered}$ (5)
where $i,j=1,2$ [7, 8]. The super-Poisson bracket between any two variable $A$ and $B$ is defined as
$\{A,B\}=\frac{\partial_{r}A}{\partial q}\frac{\partial_{l}B}{\partial p}-(-1)^{\eta_{A}\eta_{B}}\frac{\partial_{r}B}{\partial q}\frac{\partial_{l}A}{\partial p},$ (6)
where $\eta_{A}$($\eta_{B}$) denote the number of fermions, called the ghost number, in A(B) and the subscript $r$ and $l$ denotes the right and left derivatives, respectively [10].
According to the BFV formalism, the nilpotent BRST charge $(Q)$ is constructed using constraints as follows
@@
$H_{m}$ $=$ $H_{c}+\bar{{\cal{P}}_{a}}V^{a}_{b}{\cal{C}}^{b},$ (9)
$=$ $\frac{1}{2}p_{r}^{2}+\frac{1}{2r^{2}}p_{\theta}^{2}+\frac{1}{2}p_{z}^{2}+\zeta(gp_{\theta}+p_{z})+V(r)+{\cal\bar{P}}_{2}\,{\cal{C}}^{1},$
where we assume that the constraints fulfill the algebra: $\{H_{c},\Omega_{b}\}=\Omega_{a}V^{a}_{b}$. At this juncture, it is worth mentioning that the BRST charge $(Q)$ is conserved and fermionic in nature. As it is straightforward to verify that [11]
$\{Q,\;H_{m}\}=0,\qquad Q^{2}=\{Q,\;Q\}=0.$ (10)
$\{\{\Phi,\;Q\},\;Q\}=0.$ (11)
Now, the effective action in the extended phase space can be written as
$S_{eff}=\int dt\,\Big{[}p_{r}\,\dot{r}+p_{\theta}\,\dot{\theta}+p_{z}\,\dot{z}+p_{\zeta}\dot{\zeta}+B_{1}\,\dot{N^{1}}+B_{2}\,\dot{N^{2}}+{\cal\bar{P}}_{1}\,{\cal{\dot{C}}}^{1}+{\cal\bar{P}}_{2}\,{\cal{\dot{C}}}^{2}+{\cal\bar{C}}_{1}\,{\cal{\dot{P}}}^{1}+{\cal\bar{C}}_{2}\,{\cal{\dot{P}}}^{2}-H_{total}\Big{]},$ (12)
where $H_{total}=H_{m}+\{Q,\;\Phi\}$. We replace $\chi^{1}$ with $\chi^{1}+\dot{N}^{1}$ and subsequently omit the terms $\int dt\,(B_{1}\,\dot{N^{1}}\,+\,{\cal\bar{C}}_{1}\,{\cal{\dot{P}}}^{1}\,)\,$$\{Q,\int dt\,{\cal\bar{C}}_{1}\,\dot{N}^{1}\}$ in the above Eq. (12). Thus, the effective action takes the following form:
$S_{eff}=\int dt\,\Big{[}p_{r}\,\dot{r}+p_{\theta}\,\dot{\theta}+p_{z}\,\dot{z}+p_{\zeta}\dot{\zeta}+B_{2}\,\dot{N^{2}}+{\cal\bar{P}}_{1}\,{\cal{\dot{C}}}^{1}+{\cal\bar{P}}_{2}\,{\cal{\dot{C}}}^{2}++{\cal\bar{C}}_{2}\,{\cal{\dot{P}}}^{2}-H_{total}\Big{]}.$ (13)
@@
$S_{eff}$ $=$ $\int dt\,\Big{[}p_{r}\,\dot{r}+p_{\theta}\,\dot{\theta}+p_{z}\,\dot{z}+B\,\dot{N}+{\cal\bar{P}}\,{\cal{\dot{C}}}+{\cal\bar{C}}\,{\cal{\dot{P}}}-\frac{1}{2}\,p_{r}^{2}-\frac{1}{2r^{2}}\,p_{\theta}^{2}-\frac{1}{2}\,p_{z}^{2}-V(r)$ (14)
- $(gp_{\theta}+p_{z})\,N-{\cal\bar{C}}\,{\cal{C}}-B\Big{(}z-\frac{1}{2}\,B\Big{)}-{\cal\bar{P}}\,{\cal{P}}\Big{]},$
@@
$p_{r}=\dot{r},\qquad p_{\theta}=r^{2}\,(\dot{\theta}-gN),\qquad p_{z}=\dot{z}-N,\qquad{\cal\bar{P}}=-{\dot{\bar{\cal{C}}}},\qquad{\cal{P}}=\dot{\cal{C}}.$ (15)
Finally, substituting Eqs. (15) into the action (14) and identifying $N$ with $\zeta$ (cf. (2)), we obtain the covariant effective action
$S_{eff}=\int\text{d}t\,\Big{[}\frac{1}{2}\,\dot{r}^{2}+\frac{1}{2}\,r^{2}\big{(}\dot{\theta}-g\zeta\big{)}^{2}+\frac{1}{2}\big{(}\dot{z}-\zeta\big{)}^{2}-V(r)+B\Big{(}\dot{\zeta}-z+\frac{1}{2}\,B\Big{)}-{\dot{\bar{\cal{C}}}}\,{\dot{\cal{C}}}-{\cal{\bar{C}}}\,{\cal{C}}\Big{]}.$ (16)
This $(S_{eff})$ is invariant under the following off-shell nilpotent BRST symmetry transformations $(\delta_{b})$
$\begin{split}{\delta}_{b}r=0,\quad{\delta}_{b}\theta=g{\cal{C}},\quad{\delta}_{b}z={\cal{C}},\quad{\delta}_{b}\zeta=\dot{\cal{C}},\quad{\delta}_{b}{\cal{C}}=0,\quad{\delta}_{b}{\cal\bar{C}}=-B,\quad{\delta}_{b}B=0,\end{split}$ (17)
as $\delta_{b}L_{eff}=\frac{d}{dt}(B\dot{\cal{C}})$.
@@
$S^{\prime}_{eff}=\int\text{d}t\,\Big{[}\frac{1}{2}\,\dot{r}^{2}+\frac{1}{2}\,r^{2}\big{(}\dot{\theta}-g\zeta\big{)}^{2}+\frac{1}{2}\big{(}\dot{z}-\zeta\big{)}^{2}-V(r)+B\Big{(}\dot{\zeta}-z+\frac{1}{2}\,B\Big{)}-{\dot{\bar{\cal{P}}}}\,{\dot{\cal{P}}}-{\cal\bar{P}}\,{\cal{P}}\Big{]}.$ (18)
@@
3 FLPR model in Cartesian coordinates: Constraint structure

### v2

# Batalin-Fradkin-Vilkovisky Quantization and Symmetries of FLPR model

@@
Abstract: We quantize the Friedberg-Lee-Pang-Ren (FLPR) model within the framework of Batalin-Fradkin-Vilkovisky (BFV) formalism. We construct the nilpotent Becchi-Rouet-Stora-Tyutin (BRST) charges using constraints and the fermionic gauge-fixing function by means of admissible gauge conditions. We also derive the BRST invariant effective action (and corresponding symmetries) of the model in both polar and Cartesian coordinates. We demonstrate that the physical states of the system are annihilated by the first-class constraints which is consistent with the Dirac formalism. Moreover, we establish the finite field-dependent BRST (FFBRST) symmetries of the FLPR model. We exhibit the interlink between the BFV-BRST gauge-fixed action and the classical gauge invariant action using FFBRST formulation.
PACS: 11.15.-q, 11.10.Ef, 11.30.-j
Keywords: FLPR model; BFV formalism; BRST and FFBRST symmetries
@@ 1 Introduction
Constrained theories are the cornerstone of modern theoretical physics, as they include gauge theories, gravity theories, super-symmetric theories etc. However, the conventional methods of quantization are not directly applicable to such theories. To this end, Dirac (and independently Anderson &\& Bergmann) developed a quantization technique - popularly known as the Dirac formalism - to deal with such constrained systems [1, 2]. In Dirac formalism, constraints are categorized into primary, secondary, tertiary etc. and further classified into first-class and second-class. Later, Faddeev and Jackiw came up with a formalism where all the constraints are treated on equal footing without any classification [3]. The Faddeev-Jackiw (FJ) formalism is a geometrically motivated approach and based on the symplectic structure of the phase space [4]. This formalism has gained a lot of interest recently and being used widely (see e.g. [5, 6] and references therein).
@@
On the other hand, to quantize such constrained systems the BRST-BFV formalism [7, 8, 9, 10] is one of the most sought after methods. It provides a systematic Hamiltonian framework to quantize a theory endowed with first-class constraints and also preserves the BRST symmetries of the system. The BFV formalism has also been generalized to include second-class theories by employing a conversion technique called the BFT formalism [11]. These methods have been explored in a variety of classical and field theoretical models such as Chiral Schwinger model, Proca theory, QED and $(2+1)$-dimensional Hořava theory are the few names [12, 13, 14, 15].
The generalization of the BRST symmetries by making the anti-commuting infinitesimal transformation parameter finite and field dependent is known as the finite field dependent BRST (FFBRST) transformations, introduced in [16]. The effective action is invariant under such transformations; however the field dependent nature of transformation parameter disrupts the symmetry of the path integral measure. The non-trivial change in the Jacobian of the path integral measure leads to the emergence of interesting and novel results. Consequently, with a judicious choice of the finite field dependent parameter, FFBRST transformations relate the generating functional corresponding to two different actions [17, 18]. This generalized BRST (FFBRST) symmetry transformation has found numerous applications in systems such as Yang-Mills theory, Bagger-Lambert Gustavsson (BLG) theory in $N=1$ superspace, ABJM theory in $N=1$ superspace, spin-s conformal field theories, fourth-order Pais-Uhlenbeck oscillator (PUO), Christ-Lee model [19, 20, 21, 22, 23, 24, 25, 26, 27, 28] etc. Nevertheless, some basic and intriguing models, such as the FLPR model, have yet to be explored using FFBRST transformations.
@@
The FLPR model is a soluble gauge model that describes the motion of a single non-relativistic particle of unit mass governed by a central potential [29]. It is a $(0+1)$-dimensional quantum mechanical model with the characteristics of Gribov ambiguity and inspired by the dimensional reduction of Yang-Mills theory [30]. In order to avoid the gauge-fixing procedure this model is being explored using a physical projector operator [31]. On the other hand, the BRST invariant path integral is also formulated by summing over all Gribov-type copies [32]. The FLPR model is being studied using an admissible gauge-fixing condition within the framework of FJ formalism [33]. Whereas the existence of some novel off-shell nilpotent and absolutely anti-commuting symmetries of the FLPR model is established in [30]. This is an interesting model which is also being investigated by means of the supervariable approach (cf. [34, 36, 35] for details). Recently, a consistent quantization of FLPR model by avoiding the Gribov copies with corresponding BRST related symmetries is performed in [37].
@@
The prime motive of our present endeavor is to quantize the FLPR model using an admissible gauge-fixing condition, in both polar and Cartesian coordinates, by employing BFV formalism. We intend to construct the nilpotent and conserved BRST charge and subsequently derive the BRST invariant gauge-fixed effective action and corresponding symmetries. Additionally, we aim to explore physicality criteria with respect to the conserved BRST charges and compare $\grave{a}-vis$ Dirac-quantization conditions. Finally, we wish to establish the FFBRST symmetries and thereby connect the gauge-fixed effective action and classical gauge invariant action of the FLPR model.
@@
The content of the present manuscript is organized as follows. In the second section, we briefly review the BFV formalism. We provide the constraint structure of the FLPR model, in polar coordinates, using Dirac’s prescription and carried out BFV quantization in the third section. Our subsequent section deals with the constraint analysis and BFV quantization of the FLPR model in Cartesian coordinates. The conserved and nilpotent BRST charges and the physicality criteria are described in Section 5. Our section 6 deals with the FFBRST formalism of the FLPR model in polar coordinates. Finally, we provide some concluding remarks.
2 BFV Formalism
BFV formalism provides a Hamiltonian framework for quantization of gauge theories by enlarging the phase space with ghost variables and preserves the BRST symmetry.
Consider a Hamiltonian $H_{c}(q_{i},p_{i})$ defined in a $2n$-dimensional phase space of canonical variables $q_{i},\,p_{i}$ (i=1,2,,n)(i=1,2,...,n), where $q_{i}$ are generalized coordinates and $p_{i}$ are generalized momenta. We assume that the constraints $\Omega_{a}(q_{i},p_{i})$ (a=1,2,,ma=1,2,...,m) are irreducible and adhere to the following constraint algebra[12]
$\big[\Omega_{a},\,\Omega_{b}\big]=i\,\Omega_{c}U_{ab}^{c},\qquad\big[H_{c},\,\Omega_{b}\big]=i\,\Omega_{a}V_{b}^{a},$ (1)
where $U_{ab}^{c}$ and $V_{b}^{a}$ are structure coefficients. We introduce additional conditions, known as gauge-fixing conditions $\Psi^{a}(q^{i},p_{i})\approx 0$, for the purpose of isolating the physical variables in the theory. The condition $\Psi^{a}$ defines a surface that cuts through every gauge-orbit once and only once. It chooses a single representative from each gauge orbit and eliminates redundant gauge degrees of freedom. Now, we obtain 2(nm)2(n-m)-dimensional phase space defined by the physical canonical variables, say $q^{*}$ and $p^{*}$. The theory of described dynamical system depends only on the canonical variables $q^{*}$ and $p^{*}$. Thus, the generating functional is given by [12]
${\cal{Z}}=\int\big[dq^{*}dp^{*}\big]\exp\Bigg[i\int dx\Big(p^{*}\dot{q}^{*}-H_{phys}(q^{*},p^{*})\Big)\Bigg],$ (2)
here, $H_{phys}$ represents the Hamiltonian defined in terms of the physical variables in the theory. The action containing constraints $\Omega_{a}\approx 0$ and the gauge-fixing condition $\Psi^{a}(q^{i},p_{i})\approx 0$ can be written as
$S=\int dt\,\Big(p_{i}\dot{q}^{i}-H_{c}-\lambda^{a}\Omega_{a}+\pi_{a}\Psi^{a}\Big),$ (3)
where $\lambda^{a}$ and $\pi_{a}$ are Lagrange multiplier canonically conjugate to each other. The gauge-fixing condition generally takes the following form
$\Psi^{a}=\dot{\lambda}^{a}+\chi^{a}(q^{i},p_{i},\lambda^{a}),$ (4)
where $\chi^{a}$ are arbitrary functions. Here, we observe that $\lambda^{a}$ become dynamically evolving and $\pi_{a}$ act as their conjugate momenta. This leads to the canonical formalism in an extended phase space.
In order to make use of the BFV formalism, we extend the phase space by introducing two sets of canonical ghost and anti-ghost variables $\big({\cal{C}}^{a},{\cal\bar{P}}_{a}\big)$ and $\big({\cal{P}}^{a},{\cal\bar{C}}_{a}\big)$ along with the auxiliary variables $\big({N}^{a},{B}_{a}\big)$ for each constraint in the theory, such that they satisfy the following super-Poisson bracket algebra [9, 10]
$\begin{gathered}\{{\cal{C}}^{a},{\cal\bar{P}}_{b}\}=i\,\delta_{b}^{a},\quad\{{\cal{P}}^{a},{\cal\bar{C}}_{b}\}=i\,\delta_{b}^{a},\quad\big[{N}^{a},{B}_{b}\big]=i\,\delta_{b}^{a}.\end{gathered}$ (5)
The super-Poisson bracket between any two variables $A$ and $B$ is defined as
$\{A,B\}=i\,\Bigg(\frac{\partial_{r}A}{\partial q}\frac{\partial_{l}B}{\partial p}-(-1)^{\eta_{A}\eta_{B}}\frac{\partial_{r}B}{\partial q}\frac{\partial_{l}A}{\partial p}\Bigg),$ (6)
where $\eta_{A}$($\eta_{B}$) equals 0 or $1$ depending on whether $A(B)$ is a boson or fermion, respectively and the subscript $r$ and $l$ denotes the right and left derivatives, respectively [12]. It is important to note that $\big[A,B\big]=i\,\{A,B\}$. A non-trivial property of Eq. (6) is that if $A$ is a fermion, then [10]
$\{A,A\}\neq 0,\qquad\big\{\{A,\,A\},\,A\big\}=0.$ (7)
In the extended phase space, the nilpotent BRST charge $(Q)$ takes the following form:
$Q={\cal{C}}^{a}\Omega_{a}\,+\,{\cal{P}}^{a}B_{a}.$ (8)
The gauge-fixing function $(\Phi)$ can be constructed as
$\Phi={\cal\bar{C}}_{a}\,\chi^{a}+{\cal\bar{P}}_{a}\,N^{a},$ (9)
where $\chi^{a}$ are the gauge-fixing conditions of the theory. The minimal Hamiltonian $(H_{m})$ is defined by
$H_{m}=H_{c}+\bar{{\cal{P}}_{a}}V^{a}_{b}{\cal{C}}^{b}.$ (10)
At this juncture, it is worth mentioning that the BRST charge $(Q)$ is conserved and fermionic in nature. As it is straightforward to verify that [13]
$\big[Q,\;H_{m}\big]=0,\qquad Q^{2}=\{Q,\;Q\}=0.$ (11)
$\{\{\Phi,\;Q\},\;Q\}=0.$ (12)
The quantum theory is defined by the generating functional in the extended phase space as
${\cal{Z}}=\int\big[{\cal{D}}\phi\big]e^{iS_{eff}},$ (13)
where $\big[{\cal{D}}\phi\big]$= $dq^{i}\,dp_{i}]\,[d{\cal{C}}^{a}\,d{\cal\bar{P}}_{a}]\,[d{\cal{P}}^{a}\,d{\cal\bar{C}}_{a}]\,[d{N}^{a}\,d{B}_{a}]$ and effective action $(S_{eff})$ is given by
$S_{eff}=\int dt\,\Big[p_{i}\,\dot{q}^{i}+B_{a}\,\dot{N^{a}}+{\cal\bar{P}}_{a}\,{\cal{\dot{C}}}^{a}+{\cal\bar{C}}_{a}\,{\cal{\dot{P}}}^{a}-H_{total}\Big],$ (14)
where $H_{total}=H_{m}-i\,\{\Phi,\;Q\}$.
3 FLPR model in polar coordinates
@@
where $\dot{r}$, $\dot{\theta}$ and $\dot{z}$ are generalized velocities corresponding to the generalized coordinates $r$, $\theta$ and $z$, respectively. Here, $\zeta$ represents a gauge variable and $g$ is coupling constant.
3.1 Constraint structure
The canonical momenta corresponding to the variables $r$, $\theta$, $z$ and $\zeta$ are given, respectively, as
$p_{r}=\dot{r},\quad p_{\theta}=r^{2}(\dot{\theta}-g\zeta),\quad p_{z}=\dot{z}-\zeta,\quad p_{\zeta}=0.$ (16)
@@ 3.1 Constraint structure
${H}_{c}=\frac{p_{r}^{2}}{2}+\frac{p_{\theta}^{2}}{2r^{2}}+\frac{p_{z}^{2}}{2}+\zeta(gp_{\theta}+p_{z})+V(r).$ (17)
@@
There are no further constraints in the theory. Thus, the theory possess only two constraints, which are first-class in nature (cf. [33] for details).
@@ 3.2 BFV quantization
We introduce two sets of canonical ghost and anti-ghost variables $\big({\cal{C}}^{a},{\cal\bar{P}}_{a}\big)$ and $\big({\cal{P}}^{a},{\cal\bar{C}}_{a}\big)$ along with the auxiliary variables $\big({N}^{a},{B}_{a}\big)$, where $a=1,2.$ According to the BFV formalism, the nilpotent BRST charge $(Q)$ is constructed using constraints as follows
@@
$H_{m}=\frac{p_{r}^{2}}{2}+\frac{p_{\theta}^{2}}{2r^{2}}+\frac{p_{z}^{2}}{2}+\zeta\,(gp_{\theta}+p_{z})+V(r)+{\cal\bar{P}}_{2}\,{\cal{C}}^{1}.$ (21)
Now, the generating functional in the extended phase space can be written as
${\cal{Z}}=\int{\cal{D}}\phi\,\exp\Big(i\,\int dt\,\Big[p_{r}\,\dot{r}+p_{\theta}\,\dot{\theta}+p_{z}\,\dot{z}+p_{\zeta}\dot{\zeta}+B_{a}\,\dot{N^{a}}+{\cal\bar{P}}_{a}\,{\cal{\dot{C}}}^{a}+{\cal\bar{C}}_{a}\,{\cal{\dot{P}}}^{a}-H_{total}\Big],$ (22)
where
${\cal{D}}\phi={\cal{D}}\big(r,p_{r},\theta,p_{\theta},z,p_{z},\zeta,p_{\zeta},N_{a},B_{a},{\cal{C}}^{a},\bar{{\cal{P}}}_{a},{\cal{P}}^{a},\bar{{\cal{C}}}_{a}\big)$
and $H_{total}=H_{m}-i\,\{\Phi,\;Q\}$. We replace $\chi^{1}$ with $\chi^{1}+\dot{N}^{1}$ and subsequently omit the terms $\int dt\,(B_{1}\,\dot{N^{1}}\,+\,{\cal\bar{C}}_{1}\,{\cal{\dot{P}}}^{1}\,)\,$$\{Q,\int dt\,{\cal\bar{C}}_{1}\,\dot{N}^{1}\}$ in the above Eq. (22). Thus, the effective action takes the following form:
$S_{eff}=\int dt\,\Big[p_{r}\,\dot{r}+p_{\theta}\,\dot{\theta}+p_{z}\,\dot{z}+p_{\zeta}\dot{\zeta}+B_{2}\,\dot{N^{2}}+{\cal\bar{P}}_{1}\,{\cal{\dot{C}}}^{1}+{\cal\bar{P}}_{2}\,{\cal{\dot{C}}}^{2}+{\cal\bar{C}}_{2}\,{\cal{\dot{P}}}^{2}-H_{total}\Big].$ (23)
@@
$S_{eff}$ $=$ $\int dt\,\Big[p_{r}\,\dot{r}+p_{\theta}\,\dot{\theta}+p_{z}\,\dot{z}+B\,\dot{N}+{\cal\bar{P}}\,{\cal{\dot{C}}}+{\cal\bar{C}}\,{\cal{\dot{P}}}-\frac{1}{2}\,p_{r}^{2}-\frac{1}{2r^{2}}\,p_{\theta}^{2}-\frac{1}{2}\,p_{z}^{2}$ (24)
- $V(r)-(gp_{\theta}+p_{z})\,N-{\cal\bar{C}}\,{\cal{C}}-B\Big(z-\frac{1}{2}\,B\Big)-{\cal\bar{P}}\,{\cal{P}}\Big],$
@@
$p_{r}=\dot{r},\qquad p_{\theta}=r^{2}\,(\dot{\theta}-gN),\qquad p_{z}=\dot{z}-N,$
${\cal\bar{P}}=-{\dot{\bar{\cal{C}}}},\qquad{\cal{P}}=\dot{\cal{C}}.$ (25)
Finally, substituting Eqs. (3.2) into the action (24) and identifying $N$ with $\zeta$ (cf. (16)), we obtain the covariant effective action as
$S_{eff}$ $=$ $\int dt\,\Big[\frac{1}{2}\,\dot{r}^{2}+\frac{1}{2}\,r^{2}\big(\dot{\theta}-g\zeta\big)^{2}+\frac{1}{2}\big(\dot{z}-\zeta\big)^{2}-V(r)$ (26)
$+$ $B\Big(\dot{\zeta}-z+\frac{1}{2}\,B\Big)-{\dot{\bar{\cal{C}}}}\,{\dot{\cal{C}}}-{\cal{\bar{C}}}\,{\cal{C}}\Big].$
This $(S_{eff})$ is invariant under the following off-shell nilpotent BRST symmetry transformations $(s_{b})$
$s_{b}r=0,\quad s_{b}\theta=g{\cal{C}},\quad s_{b}z={\cal{C}},\quad s_{b}\zeta=\dot{\cal{C}},\quad s_{b}{\cal{C}}=0,\quad s_{b}{\cal\bar{C}}=-B,\quad s_{b}B=0,$ (27)
as $s_{b}L_{eff}=\frac{d}{dt}(B\dot{\cal{C}})$.
@@
$S^{\prime}_{eff}$ $=$ $\int dt\,\Big[\frac{1}{2}\,\dot{r}^{2}+\frac{1}{2}\,r^{2}\big(\dot{\theta}-g\zeta\big)^{2}+\frac{1}{2}\big(\dot{z}-\zeta\big)^{2}-V(r)$ (28)
$+$ $B\Big(\dot{\zeta}-z+\frac{1}{2}\,B\Big)-{\dot{\bar{\cal{P}}}}\,{\dot{\cal{P}}}-{\cal\bar{P}}\,{\cal{P}}\Big].$
@@
4 FLPR model in Cartesian coordinates